OpenStudy (anonymous):

The function h(t) = -16t squared + 650 models the height in feet of rock is dropped from a 650 foot building. How long will it take to hit the ground. (round to the nearest hundredth second)

OpenStudy (loser66):

Let h(t) =0, solve for t

OpenStudy (anonymous):

so substact the 650 to both sides then divide by -16

OpenStudy (loser66):

yup

OpenStudy (anonymous):

do u do anything with the square

OpenStudy (loser66):

sure, take square root of the result above to get the answer.

OpenStudy (anonymous):

i got a decimal when dividing

OpenStudy (loser66):

@perl at the first glance I thought your avatar is a bunch of flower. hahaha

OpenStudy (perl):

:)

OpenStudy (loser66):

@davidellis17 so?? it doesn't matter

OpenStudy (loser66):

@perl I know you are a man, hence it 's weird if the avatar is flower.

OpenStudy (anonymous):

ok so once your done dividing to get your answer you then take the square of tht answer?

OpenStudy (loser66):

\[\sqrt{\dfrac{-650}{-16}}\]

OpenStudy (anonymous):

ok im good on tht one but i have more

OpenStudy (anonymous):

write the polynomial 2x(x-2)(x+1) in standard form. then classify it by degree and number of terms.

OpenStudy (loser66):

just distribute the terms, like open the brackets. dat sit

OpenStudy (anonymous):

ohh gotcha

OpenStudy (loser66):

:)

OpenStudy (anonymous):

\[^{3}\sqrt{a ^{16}}b ^{19}\] simplify